# MAnycast Reloaded: a Tool for an Open, Fast, Responsible and Efficient Daily Anycast Census

### v1

@@ 5.5.2. Reducing probing rate
A key strength of MAnycastR is a low probing cost and hence a responsible measurement of IP anycast. To illustrate this, we test another feature of MAnycastR. We can configure MAnycastR to go over the hitlist at a specific rate, while maintaining one-second intervals between probes from each Worker. To validate if we can maintain accuracy at a reduced probing rate, we perform a MAnycastR census using a probing rate of 1/8th the ‘normal’ rate of our daily census. Even at this much-reduced rate, MAnycastR detects the same number of anycast targets.
@@ 5.6. Anycast prefix size
As mentioned in §4, our methodology scans at a /24 (and /48 for IPv6) granularity as it is the smallest prefix size propagated by BGP. However, BGP announcements seen by route collectors are often less specific. To assess whether such less specific prefixes are anycast in their entirety, we look at all announced IPv4 prefixes in which we detect anycast.
@@
This comparison is done using data from April 27, 2024, where our census finds 12,046 GCD-confirmed anycast /24-prefixes. Using CAIDA’s prefix2as dataset (CAIDA, 2024) we find those /24s to be part of 4,184 BGP announced prefixes. Out of these BGP prefixes 3,827 are entirely anycast (i.e., each contained /24-prefix is anycast). For 70 prefixes there is uncertainty as they contain at least one /24 that is unresponsive for the measurement. The remaining 287 announced prefixes contain unicast /24-prefixes according to our measurement.
@@
This shows scanning at /24 granularity is necessary to avoid overestimating anycast. However, routing may yet be different within such prefixes (Schomp and Al-Dalky, 2020) . This limitation was evident when validating results with NTT DATA Global IP Network, a large tier-1 transit provider, who announce their address space at multiple PoPs (i.e., global BGP). However, their addresses point to single servers (i.e., unicast). The exception being 6 addresses (in a single /24) replicated at all PoPs for their public DNS resolver (i.e., anycast). The census, which probed a unicast address for this prefix, falsely classified the entire /24 as unicast. We call such cases, where a /24 contains both unicast and anycast, partial anycast.
To assess the number of partial anycast prefixes, we perform a GCD measurement at /32 granularity targeting the entire allocated IPv4 address space, totaling four billion targets. We do this using nine VPs spanning multiple continents, at a low probing rate, and over a period of ten days.
@@
While GCD measurements require a large number of VPs for accurate detection, we believe in this case a few VPs suffice for detecting partial anycast as it requires a private backbone with global reach to allow for different routing topologies within an announced prefix.
This scan revealed anycast in 13.4k /24s, of which 1,483 /24s contain partial anycast (including the aforementioned tier-1 prefix). Following this scan we updated our AT list and included a partial anycast flag in our census to avoid overestimating anycast. 1,178 of these partial anycast prefixes are consistently found in our census. We observe most partial anycast prefixes are announced by large CDNs that have global presence. For 305 partial anycast prefixes we find the prefix is entirely unicast the following day. These prefixes mostly belong to Imperva, which we observe to announce temporary anycast in our RIPE Atlas measurements and longitudinal census data. We suspect this is related to their anti-DDoS services. This motivates our decision to measure anycast longitudinally. Future work is to capture short-lived anycast by triggering measurements based on, e.g., BGP announcements observed by route collectors.
@@ 5.7. Comparing with external sources
There are also external datasets that contain information on anycast. We compare our work against two, a commercial dataset from IPInfo and a public dataset from BGPTools.
IPInfo, a commercial provider of IP address databases, shared their list of detected anycast prefixes with us (ipinfo.io, 2025).
First, for IPv6 our census detects anycast in 6.3k /48s whereas IPInfo detects anycast in 2.0k /48s. For IPv4, our census detects anycast in 13.4k /24s and IPInfo in 14.0k /24s. Overall, 12.6k /24s are found in both datasets showing high agreement. Inspecting the 0.8k prefixes found only in our census, we find most are detected in few locations and are regional. For the 1.4k prefixes found only by IPInfo, we observe 0.6k originate from Imperva and several hundreds from other CDNs offering anti-DDoS services where we suspect temporary anycast. We believe IPInfo to include larger numbers of temporary anycast as they take weekly snapshots.
BGPTools also uses an anycast-based approach similar to the first stage of MAnycastR. However, there are two key differences: 1) if BGPTools detects a single address in an announced prefix as anycast, they assume the entire prefix is anycast (BGPtools, 2024), and 2) they do not use GCD to filter out FPs. We compare BGPTools to MAnycastR for December 20, 2024. On that day, BGPTools marks 3,047 announced prefixes as anycast. Comparing this to MAnycastR, of those 3,047 prefixes our anycast-based stage marks 2,954 as anycast. Of these, 228 prefixes are only detected at two of our VPs, and our GCD stage marks these as FPs, showing BGPTools likely significantly overestimates the number of prefixes that use anycast. Our census finds 13,495 anycast /24-prefixes confirmed by GCD in total for that day. Of these, 9,739 are also covered by the BGPTools prefixes, leaving 3,756 /24-prefixes they miss.
Investigating the prefixes BGPTools classifies as anycast, we find 467 are less-specific than a /24. Comparing this to our census, we find that 60 of these prefixes contain a grand total of 8,038 unicast /24s, illustrating the risk of BGPTools’ assumption that entire announced prefixes are anycast.
Looking at IPv6, BGPTools detects 1,148 IPv6 prefixes, 1,131 of which are covered by our census while 8 are missed as they are not in our hitlist. That same day we find 6,358 anycast /48s, 1,479 of which were not found by BGPTools.
@@
5.8. Ground truth validation
Finally, we validate against public datasets and reach out to operators to obtain ground truth for our census results.
5.8.1. DNS operators
We find all IPv4 and IPv6 prefixes for the Quad9 public resolver, RIPE authoritative nameservers, and all DNS root servers correctly classified as anycast by our census. Interestingly, we observe that G-root is unresponsive to ICMP and TCP for both IPv4 and IPv6. However, MAnycastR is able to detect it using DNS over UDP probing.
5.8.2. Hypergiants
Table 6 shows the largest ASes based on the number of IPv4 /24 and IPv6 /48 anycast prefixes they originate. We find that Google Cloud is leading in IPv4 anycast followed by Cloudflare, Amazon, and Fastly. These ASes are large CDNs, often referred to as hypergiants. We color the table based on ground truth validation; green for operator-confirmed ground truth (no FPs), dark-green when fully accurate (no FPs and FNs), and orange when confirmed with public data. Due to the dominance of hypergiants, these results comprise 59% of our IPv4 and 63% of our IPv6 census.
@@ 5.8.2. Hypergiants
Table 6. Largest ASes originating anycast prefixes.
@@
Google provides an ipranges dataset (Google, 2025a) that discloses where prefix ranges are announced for Google Cloud resources. We compared our census to this dataset for January 6, 2025. Google’s dataset contains 33 globally announced IPv4 prefixes ranging from a /16 to a /24, totaling 3,581 /24-prefixes. Out of these prefixes, 3,572 are detected as anycast in our census. For the 9 missing prefixes, 8 are not in our hitlist and 1 we detect as unicast (which we confirm to be unicast using traceroute). The remaining Google Cloud prefixes we detect as anycast are not listed in the ipranges dataset.
@@
Cloudflare, the CDN that ranks first for IPv6 and second for IPv4 anycast in the census, shared ground truth data with us. Their data confirms our census is fully accurate for IPv4 anycast, where we have no FPs and no FNs. For IPv6 they shared we have no FPs, however, our census misses some IPv6 prefixes due to IPv6 hitlist limitations.
@@
Next, we evaluate the prefixes we detect as anycast for Amazon using an ip-ranges dataset (Amazon, 2025) similar to Google. The list contains 37k /24 prefixes listed as globally announced, these are cases of global announcements but not necessarily anycast as discussed previously. This list includes partitions more specific than /24, even /32 partitions. For the 1,286 /24-prefixes we detect as anycast, 1,123 are listed as being announced globally, 2 are listed as being announced in the east coast of the US, and the remaining 161 are not listed.
@@
We also reached out to Fastly whose prefixes feature prominently in our dataset and obtained ground truth data. Furthermore, they disclosed a traffic engineering implementation that uses backing anycast, where a large prefix is announced globally (anycast) and more specific prefixes within route to a single PoP (unicast), which allows for withdrawing the unicast announcement to fall back on the less-specific anycast announcement. Such traffic engineering allows for rapid redistribution of traffic, e.g., if a unicast prefix announced at a single PoP is the victim of a DDoS attack. For IPv4 Fastly confirmed all /24-prefixes we find are anycast. However, for IPv6 we have a considerable number of misclassifications. After troubleshooting, we discovered that two Ark nodes are located in ASes that miss /48 announcements for the CDN due to these ASes filtering these out or their upstream ASes not propagating the announcements. In the case of these FPs the probed addresses are unicast using a /48 with a less specific backing anycast prefix (the TE approach mentioned above). Due to these nodes being unaware of the /48, they are routed to the nearest PoP, the address is then misclassified as anycast. Filtering out these Ark nodes from our IPv6 results, we find all detected IPv6 anycast to match ground truth.
@@
These findings indicate our methodology can detect ASes that, e.g., do not propagate /48-prefixes, by performing MAnycastR measurements where we announce a specific prefix at a single site. Finally, the TE practice used to switch prefixes from unicast to anycast when a PoP is overloaded motivates the need for a daily census.
5.9. Lessons learned

### v3

@@ 5.5.2. Reducing probing rate
A key strength of LACeS is a low probing cost and hence a responsible measurement of anycast. To illustrate this, we test another feature of LACeS. We can configure LACeS to go over the hitlist at a specific rate, while maintaining one-second inter-worker spacings (i.e., time between Workers sending their probe toward the same target). To validate if we can maintain accuracy at a reduced probing rate, we perform a LACeS anycast-based measurement using a probing rate of 1/8th the ‘normal’ rate of our daily census. Even at this much-reduced rate, LACeS detects the same number of anycast candidates.
5.6. Topology impact
The previous section explored the effects of using different VPs and hosting providers, showing the coverage of the anycast-based measurement depends on the number and locations of VPs used. However, as noted in the MAnycast2 paper (Sommese et al., 2020), the number and diversity of upstreams at each location may have a large impact on coverage as well. To further explore the effects of topological connectivity we change routing at each location using BGP communities supported by our upstream. These measurements were performed in May. ’25. In particular, we use Vultr’s Do not announce to IXP peers and Announce to IXP route servers only (Vultr, 2025a). The former means our anycast prefix is only propagated to transit ASes (Transits-only) and the latter means our prefix is propagated only to IXP peers (IXPs-only), thereby changing the path taken by reply traffic. We provide a Venn diagram of the ACs found using the various announcements in Figure 8.
Figure 8. ACs found using different routing policies.
Transits-only found the most ACs (36,582) followed by IXPs-only (27,345) and the unmodified announcements found the fewest (25,899). We suspect the Transit-only community found more ACs as transit ASes may have equal-cost paths to multiple PoPs, where replies from unicast targets reach multiple VPs (e.g., due to ECMP).
Filtering on prefixes that are GCD-confirmed, we find that IXP-only found the most (13,033) followed by the unmodified announcement (13,014) and Transits-only found the fewest (12,713). For these prefixes we find the anycast-based approach failed to detect them as candidate anycast using certain communities. Surprisingly, the Transits-only community with the most ACs had the fewest GCD-confirmed prefixes. We speculate that some transit ASes may have a shortest or preferred path to a single VP causing anycast targets to reply to fewer VPs. For instance, 1.1.1.0/24 has replies reach 31 VPs for both the unmodified and IXPs-only announcements whereas it reaches only 16 VPs for the Transits-only announcement.
These results give an indication as to how different routing topologies impact results. Furthermore, it suggests that transit AS routing policies may have negative effects on the accuracy of the anycast-based approach.
@@ 5.7. Anycast prefix size
BGP prefixes – As mentioned in §4, our methodology scans at a /24 (and /48 for IPv6) granularity as it is the smallest prefix size propagated by BGP. However, BGP announcements seen by route collectors are often less specific. To assess whether such less specific prefixes are anycast in their entirety, we look at all announced IPv4 prefixes in which we detect anycast. This comparison is done using data from Apr. ’24, where our census found 12,046 GCD-confirmed anycast /24-prefixes ($\mathscr{G}$). Using CAIDA’s prefix2as dataset (CAIDA, 2024) we find those /24s to be part of 4,184 BGP announced prefixes. Out of these BGP prefixes, 3,827 (91%) are entirely anycast (i.e., each contained /24-prefix is anycast). For 70 prefixes (2%) there is uncertainty as they contain at least one /24 that is unresponsive for the measurement. The remaining 287 BGP prefixes (7%) contain unicast /24-prefixes according to our measurement. This motivates our decision to scan at /24 granularity, rather than scanning at BGP announcement granularity which leads to overestimating anycast.
@@
Partial anycast – However, routing may yet be different within such prefixes (Schomp and Al-Dalky, 2020) . This limitation was evident when examining inferences made with NTT DATA Global IP Network, a large tier-1 transit provider, who announce their address space at multiple PoPs (i.e., global BGP). However, their addresses point to single servers (i.e., unicast). The exception being 6 addresses (in a single /24) replicated at all PoPs for their public DNS resolver (i.e., anycast). The census, which probed a unicast address for this prefix, classified the entire /24 as unicast. We call such cases, where a /24 contains both unicast and anycast, partial anycast.
@@
To assess the number of partial anycast prefixes, we perform a GCD measurement at /32 granularity targeting the entire allocated IPv4 address space, totaling four billion targets. We performed this measurement in Nov. ’24 and Aug. ’25, which we refer to as GCDIPv4. For the Aug. ’25 measurement we used 13 VPs spanning multiple continents over a period of ten days. While GCD measurements typically require a large number of VPs for accurate detection, we believe in this case a few VPs suffice for detecting partial anycast as partial anycast requires a private backbone with global reach to allow for different routing topologies within an announced prefix.
This scan revealed anycast in 12.4k /24s, for which we inferred partial anycast in 1,025 (8%) /24s (including the aforementioned tier-1 prefix). We include a partial anycast flag in our census to avoid overestimating anycast. We observe most partial anycast prefixes are announced by large CDNs that have global presence.
Surprisingly, we found more (1,483) partial anycast prefixes in Nov. ’24 of which 793 were not found in Aug. ’25. Inversely, 335 partial prefixes in Aug. ’25 were not found in Nov. ’24. Investigating these non-intersecting prefixes, we observe they are mostly seen as entirely unicast within a short period. Additionally, they are largely originated by Imperva which we suspect to be temporary anycast as observed in our RIPE Atlas measurements.
This analysis suggests partial anycast may be short-lived in many cases. While interesting to study, the high measurement cost (a full GCD of the IPv4 address space) and impact of such measurements make a longitudinal study of this phenomenon undesirable. Our current plan is to repeat GCDIPv4 on an annual basis to capture the fact that such dynamics exist while accepting the limitation that we cannot capture the dynamics of this phenomenon in more detail
@@ 5.8. Comparing with external sources
There are also external datasets that contain information on anycast. We compare our work against two, a commercial dataset from IPInfo and a public dataset from BGPTools using census data from Dec. ’24. As our census has high confidence in GCD-confirmed prefixes ($\mathscr{G}$) we focus the comparison on these results.
IPInfo, a commercial provider of IP address databases, shared their list of detected anycast prefixes with us (ipinfo.io, 2025). First, for IPv6 our census detects anycast in 6.3k /48s whereas IPInfo detects anycast in 2.0k /48s. For IPv4, our census detects anycast in 13.4k /24s and IPInfo in 14.0k /24s. Overall, 12.6k /24s are found in both datasets showing high agreement. Inspecting the 0.8k prefixes found only in our census, we find most are detected in few locations and are regional. For the 1.4k prefixes found only by IPInfo, we observe 0.6k originate from Imperva and several hundreds from other CDNs offering anti-DDoS services where we suspect temporary anycast. We believe IPInfo to include larger numbers of temporary anycast as they accumulate anycast prefix using weekly snapshots.
BGPTools uses an anycast-based approach similar to the first stage of LACeS. However, there are two key differences: 1) if BGPTools detects a single address in an announced prefix as anycast, they classify the entire prefix is anycast (BGPtools, 2024), and 2) BGPTools do not use GCD to filter out FPs. BGPTools marks 3,047 announced prefixes as anycast, of those 3,047 prefixes our anycast-based stage marks 2,954 (97%) as anycast. Of these, 228 (8%) prefixes are detected at two of our VPs, and our GCD stage marks these as FPs, suggesting BGPTools overestimates the number of prefixes that use anycast. Our census finds 13,495 anycast /24-prefixes confirmed by GCD ($\mathscr{G}$) in total for that day. Of these, 9,739 (72%) are also covered by the BGPTools prefixes, leaving 3,756 (28%) /24-prefixes they miss.
Investigating the prefixes BGPTools classifies as anycast, we find 467 (15%) are less-specific than a /24. Comparing this to our census, we find 60 of these prefixes contain a grand total of 8,038 unicast /24s, illustrating the risk of classifying entire announced prefixes as anycast.
Looking at IPv6, BGPTools detects 1,148 IPv6 prefixes, 1,131 of which are covered by our census while 8 are missed as they are not in our hitlist. That same day we find 6,358 anycast /48s, 1,479 (23%) of which were not found by BGPTools.
@@
6. Ground truth validation
Finally, we validate our GCD-confirmed results ($\mathscr{G}$) against public datasets and reach out to operators to obtain ground truth.
@@ 6. Ground truth validation
Table 5. Largest ASes by number of anycast /24-prefixes.
DNS operators – We find all IPv4 and IPv6 prefixes for the Quad9 public resolver, RIPE authoritative nameservers, and all DNS root servers correctly classified as anycast by our census. Interestingly, we observe that G-root is unresponsive to ICMP and TCP for both IPv4 and IPv6. However, LACeS is able to detect it using DNS.
Hypergiants – Table 5 shows the largest ASes based on the number of IPv4 /24 and IPv6 /48 anycast prefixes they originate. We find that Google Cloud is leading in IPv4 anycast followed by Cloudflare, Amazon, and Fastly. These ASes are large CDNs, often referred to as hypergiants. We color the table based on ground truth validation; green for operator-confirmed ground truth (no FPs), dark-green when fully accurate (no FPs or FNs), and orange when confirmed with public data. Due to the dominance of hypergiants, these results comprise 59% of our IPv4 and 63% of our IPv6 census.
@@
Google provides an ipranges dataset (Google, 2025a) that discloses where prefix ranges are announced for Google Cloud resources. We compared our census to this dataset for January 6, 2025. Google’s dataset contains 33 globally announced IPv4 prefixes ranging from a /16 to a /24, totaling 3,581 /24-prefixes. Out of these prefixes, 3,572 are detected as anycast in our census. For the 9 missing prefixes, 8 are not in our hitlist and 1 we detect as unicast (which we confirm to be unicast using traceroute). The 55 remaining Google Cloud prefixes we detect as anycast are not listed in the ipranges dataset.
@@
Cloudflare, the CDN that ranks first for IPv6 and second for IPv4 anycast in the census, shared ground truth data with us. Their data confirms our census is fully accurate for IPv4 anycast, where we have no FPs and no FNs. For IPv6, we have no FPs, however, our census misses IPv6 prefixes due to IPv6 hitlist limitations.
@@
Next, we evaluate the prefixes we detect as anycast for Amazon using an ip-ranges dataset (Amazon, 2025) similar to Google. The list contains 37k /24 prefixes listed as globally announced, these are cases of global announcements but not necessarily anycast as discussed previously. This list includes partitions more specific than /24, even /32 partitions. For the 1,286 /24-prefixes we detect as anycast, 1,123 (87%) are listed as being announced globally, 2 are listed as being announced in the east coast of the US, and 161 (13%) are not listed.
@@
We also reached out to Fastly whose prefixes feature prominently in our dataset and obtained ground truth data. They disclosed a traffic engineering implementation that uses backing anycast, where a large prefix is announced globally (anycast) and more specific prefixes within route to a single PoP (unicast), which allows for withdrawing the unicast announcement to fall back on the less-specific anycast announcement. Such traffic engineering allows for rapid redistribution of traffic, e.g., if a unicast prefix announced at a single PoP is the victim of a DDoS attack. For IPv4 Fastly confirmed all /24-prefixes we find are anycast. However, for IPv6 we have a considerable number of FPs. After troubleshooting, we discovered that two Ark nodes are located in ASes that miss /48 announcements for the CDN due to these ASes filtering these out or their upstream ASes not propagating the announcements. In the case of these FPs the probed addresses are unicast using a /48 with a less specific backing anycast prefix (the TE approach mentioned above). Due to these nodes being unaware of the /48, they are routed to the nearest PoP, the address is then misclassified as anycast. Filtering out these Ark nodes from our IPv6 results, we find all detected IPv6 anycast to match ground truth.
@@
These results indicate our methodology detects ASes that, e.g., do not propagate /48-prefixes, by performing anycast-based measurements where we announce a specific prefix at a single site. Finally, the TE practice used to switch prefixes from unicast to anycast when a PoP is overloaded motivates the need for a daily census.
Figure 9. LACeS detection statistics by method and protocol over time for both IPv4 and IPv6.
7. Longitudinal View
This section provides a longitudinal view of LACeS, covering daily census data between Mar. ’24 and Sep. ’25 (17+ months).
Hitlist and feedback loop – The coverage of our census is dependent on the hitlists used, which we update quarterly to maintain high coverage. First, the USC/ISI hitlist (Fan and Heidemann, 2010) updates quarterly due to churn in ping-responsive hosts. However, we observe this churn mostly affects unicast hosts and has limited effects on LACeS. The change in coverage for IPv6 can be explained by the growth of the hitlist (Gasser et al., 2018).
Furthermore, prefixes found using GCDLS are fed into the feedback loop to cover rare cases of regional anycast undetectable by the anycast-based approach. We conducted the GCDLS measurement in Feb. ’24, Dec. ’24, and Aug. ’25 finding 13,684, 13,692, 13,514 /24s respectively. Comparing those found in Feb. ’24 and Aug. ’25, we find an intersection of 11,719 /24s that remained anycast. For 1,965 /24s found in Feb. ’24, we observe they are no longer anycast and 1,796 new prefixes are detected as anycast in Aug. ’25. These results highlight the need for frequent GCDLS measurements. For IPv6 the GCDLS sets had 4,891, 6,314, and 11,286 /48s respectively (on the same dates as IPv4). Unlike IPv4, we see large variations as coverage is largely limited by the IPv6 hitlists used.
Similarly, we periodically repeat GCDIPv4 sweeps to provide coverage for partial anycast and ping-responsive anycast prefixes missed by our hitlists.
Daily ACs – To demonstrate the value of daily ACs from the anycast-based approach (instead of solely relying on prefixes found using GCDLS) we compare Dec. ’24 GCDLS prefixes with three snapshots in January, May, and August 2025. We find that as time progresses our daily results deviate substantially from this one time snapshot. We find 40, 255, 399 GCD-confirmed prefixes from daily ACs on the respective snapshots that were not found with GCDLS in Dec. ’24. These results clearly show the value of daily ACs to maintain coverage (in particular for short-lived anycast).
Detection counts – Figure 9 shows the number of prefixes found using all of our probing methods for both IPv4 and IPv6. First, looking at the IPv4 statistics we observe variability for the number of ACs found using the anycast-based measurements. This is due to routing dynamics causing false positives in the anycast-based approach as explained in §5.1.3. For GCD we observe an increase in Dec. ’24 when we updated the feedback loop with GCDLS prefixes and Jan. ’25 when the number of Ark VPs increased.
Looking at IPv6, we observe growth in ACs and GCD-confirmed prefixes as our IPv6 hitlist grows. Most notably in Aug. ’25 where the number of GCD-confirmed prefixes doubled. In Apr. ’24, we see an uptick in IPv6 ACs belonging to China Unicom followed by large increases in April-June and July from Astound Broadband (a large US telco) and a peak in August from contell (a Russian retailer). We suspect these ASes experienced routing instability affecting a large number of /48s due to large IPv6 prefix sizes. Interestingly, we confirm the 20 k /48s increase in ACs from July using GCD. Closer investigation shows these belong to two /29 IPv6 prefixes announced by Astound (visible in the public routing table), which iGreedy locates to Baltimore and New York City.
Figure 10. Cumulative count of anycast /24-prefixes detected to be anycast for at most “X” days.
Persistent anycast – Figure 10 shows the cumulative counts of anycast prefixes found during the census period, based on the number of days they were detected as anycast. The first thing that stands out, on the right of the graph, are 9.7 k (60%) /24s, belonging to 433 distinct ASes, that were anycast for the entire period. Of these prefixes, 66% were confirmed via ground truth in §6; 3.2 k (33%) from Google, 2.6 k (27%) from Cloudflare and 1.2 k (12%) from Amazon.
We find 3 k (19%) prefixes observed to be anycast for only 20-80% of the measured period, 603 of these belong to Google (AS 396982) and 402 to Fastly (AS 54113) where we suspect dynamic address utilization (e.g., backing anycast as discussed in §6). For the remainder there exist multiple reasons; i) changes in LACeS coverage, ii) prefixes that became anycast during the period, iii) or prefixes that dynamically switch between unicast and anycast. Cases ii) and iii) can only be observed through using the daily ACs from the anycast-based measurement
The bottom left shows cases of temporary anycast. In total, we find 1,069 /24s that were anycast for a maximum of 10 days during our census period. Of these, 572 belong to Imperva (AS 19551) that offers content caching, and DDoS mitigation services for customers using their anycast network. Finally, we find 191 prefixes that were anycast only for a single day. Manual inspection shows these are mostly anycasted at two locations (i.e., the unicast location and an anomalous second location). We suspect these are BGP misconfiguration or hijacking events, matching observations by Cicalese et al. that hypothesized iGreedy could detect BGP hijacking. Future work is detection of such events using LACeS, including detection based on BGP data to cover short-lived events not visible with daily measurements.
Enumeration – We observe enumeration for large anycast deployments increasing longitudinally as the Ark platform expands (e.g., for Cloudflare we find $\sim$30 sites in Mar. ’24 compared to $\sim$70 sites in Sep. ’25). Furthermore, we observe deployment changes in our daily census, e.g., during our census we find additional locations for .cz matching infrastructure expansions disclosed by the operator.
Technical events – From 19th of September 2024 till 24th of December 2024, our census collected no DNS results due to issues with the tooling that was incorrectly flagging all replies as invalid. Following this issue we added an alerting system that triggers when canary checks fail or results substantially deviate from the baseline.
In Figure 9 we observe large drops in ICMPv4 ACs, these are caused by Workers losing connection with the Orchestrator resulting in anycast-based measurements with fewer VPs. We fixed this issue in Jul. ’25 by implementing automatic re-connects thus improving robustness of LACeS. We do note this issue had minor effects in GCD-confirmed prefixes thanks to the feedback loop in LACeS.
8. Discussion